I am having trouble helping my son with the below problem .
Tom's father is 3 times old as Tom. 4 years ago, he was 4 times as older. How old is Tom?
Not sure how to set this up and solve. Can you help me?

Answer by richwmiller(17219)

Turn the words into equations
f=3t
f-4=4*(t-4)
You will need an equation for each unknown
We have two unknowns (father and Tom )and we have two equations.
Now substitute 3t in for f in the 2nd equation
3t-4=4*(t-4)
and solve for t
This is nice because we want to know t (Tom's age)